ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

11. Mathematical Physics

1991 MS Classification: Bogomolny, E. B., CNRS-URA 62, Institut de Physique Nucléaire, Orsay, France Spectral Statistics

The aim of the lecture is to present an overview of statistical properties of energy eigenvalues of deterministic systems and related questions. The following subjects are shortly discussed: Distribution of energy eigenvalues for integrable and chaotic systems. Universality within standard random matrix ensembles. Distribution of zeros of number-theoretical zeta functions. Arithmetical models. Intermediate statistics. Approach to universal limit. Random polynomials.

ICM 1998 Berlin · Plenary and Invited Lectures